OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5).

OpenStudy (anonymous):

^2

OpenStudy (anonymous):

2x^2

OpenStudy (solomonzelman):

Plug in 5 for x, and find out.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x)=2x^2+2}\) \(\normalsize\color{blue}{ f(2)=2(2)^2+2~~~~~~~⇒~~~~~~~f(2)=10}\) \(\normalsize\color{blue}{ f(3)=2(3)^2+2~~~~~~~⇒~~~~~~~f(3)=20}\) \(\normalsize\color{blue}{ f(2)=2(4)^2+2~~~~~~~⇒~~~~~~~f(4)=34}\)

OpenStudy (solomonzelman):

The last one I posted is f(4) And on... So, \(\normalsize\color{blue}{ f(5)=2(5)^2+2~~~~~~~⇒~~~~~~~f(5)=??}\) you tell me

OpenStudy (anonymous):

I don't know what to do with the f(5)? I'm still confused

OpenStudy (solomonzelman):

\(\normalsize\color{darkgoldenrod}{ f(5)=2(5)^2+2}\) . You just need to simplify the `2(5)²+2` to find the ` f(5)`.

OpenStudy (anonymous):

52? @SolomonZelman

OpenStudy (solomonzelman):

yes, 52 is right !

OpenStudy (anonymous):

thank you sooo much!

OpenStudy (solomonzelman):

Anytime:)